OpenStudy (anonymous):

What is the d value in the following arithmetic sequence? -7, -2, 3, 8, …

OpenStudy (amistre64):

what does the value of d do for us in an arithmetic sequence?

OpenStudy (anonymous):

it's the common difference but that doesn't really make sense to me.. the course explains it as the common number but what does that mean?

OpenStudy (amistre64):

good, it means that in order to get from one number, to the next number, we add the same amount each time. in this case: -7 + d = -2 -2 + d = 3 3 + d = 8

OpenStudy (solomonzelman):

Do you need some explanation of this in general, like what this all about, `difference` `sequence` `arithmetic` and all this ?

OpenStudy (anonymous):

ohhh wait so you would be adding 5each time. is that the value of d?

OpenStudy (amistre64):

you would be adding 5 each time :) so yes, d=5

OpenStudy (anonymous):

okay thank you very much

OpenStudy (amistre64):

youre welcome

OpenStudy (solomonzelman):

aswwsa, can you find the 15 number in this sequence ?

OpenStudy (solomonzelman):

I'll post a brief explanation of other notations (of how this is usually written) okay ?

OpenStudy (anonymous):

that would help a lot

OpenStudy (solomonzelman):

Lets take your sequence for example \(\normalsize\color{black}{ -7, ~-2, ~3, ~8}\) the first term is `-7` the first term is `-2` the first term is `3` the first term is `8` In general though, the terms are labeled like this: \(\large\color{black}{ a_1}\) first term \(\large\color{black}{ a_2}\) second term \(\large\color{black}{ a_3}\) third term \(\large\color{black}{ a_4}\) forth term and on.....

OpenStudy (solomonzelman):

So in any sequence, how would you write the 100th term ?

OpenStudy (anonymous):

that would be al00 right?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

Would it make sense to say that to get from \(\large\color{black}{ a_1}\) to \(\large\color{black}{ a_{100} }\) you add the difference (whatever it is) 99 times ?

OpenStudy (anonymous):

you mean like add d 99 times? no

OpenStudy (solomonzelman):

YES !!!!very good :)

OpenStudy (solomonzelman):

So we can say that \(\large\color{black}{ a_{n}=a_1+d(n-1) }\) would this make sese ?

OpenStudy (solomonzelman):

*sense

OpenStudy (anonymous):

ohh yeah that makes more sense so say I'm finding the 25th term you would it end up like this.. an=-7+5(25-1) ?

OpenStudy (solomonzelman):

yes if it is \(\large\color{black}{ a_{25}=-7+5(25-1) }\) if you were to look for a 25th term in our sequence. and what would \(\large\color{black}{ a_{25} }\) be equal to, can you calculate that ?

OpenStudy (solomonzelman):

(Sorry for late reply, just finished my college application)

OpenStudy (anonymous):

it's fine. it would be 113?

OpenStudy (solomonzelman):

yes !

OpenStudy (solomonzelman):

Can you find the 50th term ?

OpenStudy (anonymous):

that would be 238

OpenStudy (solomonzelman):

yes !

OpenStudy (solomonzelman):

Now, lets say if you needed to add all 50 terms up, you would use a formula \(\LARGE\color{midnightblue}{ \rm S_n=\color{red}{\frac{1}{2}(a_1+a_n)}× n }\) Sn is 'sum of n number of terms' n is number of terms

OpenStudy (solomonzelman):

\(\LARGE\color{midnightblue}{ \rm S_{50}=\color{red}{\frac{1}{2}(-7+238)}× 50 }\)

OpenStudy (anonymous):

ok 1 more quick question before that.. how would you find the 4 arithmetic terms between -21 and -36

OpenStudy (solomonzelman):

4 arithmetic means between 21 and 36

OpenStudy (solomonzelman):

-21 and -36 •••• So, just for imagination will say -21, __ , __ , __ , __, -36 \(\normalsize\color{blue}{ -36=-21+d(5)}\) (because you should add d 5 times to get from -21 to -36)

OpenStudy (solomonzelman):

Solve for d, and then you will be able to fill in the missing terms.

OpenStudy (solomonzelman):

if you had N missing means, you would add d `N+1` times

OpenStudy (anonymous):

it would be -3 I think?

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

adding 21 to both sides, ad then dividing both sides by -3 :)

OpenStudy (solomonzelman):

So if d=-3 then the 4 missing terms are ?

OpenStudy (anonymous):

it'll be -24, -27, -30, -33

OpenStudy (solomonzelman):

yes !

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ S_n=\underline{\frac{1}{2} (a_1+a_n)} ×n }\) Sn is `sum of n number of terms` n is ` number of terms ` I underlined the average term, makes sense why this is the average term in ANY sequence ?

OpenStudy (anonymous):

yes I think that makes sense. another question.. sorry I have so many.. how would you find S25 for -7+ -2 + 3 + 8 + ... ?

OpenStudy (solomonzelman):

Use the formula I posted

OpenStudy (anonymous):

ohhh would you just find the value of the 25th term and add that to the first the take half of that? what would you multiply it by then, what is n?

OpenStudy (solomonzelman):

a25 = a1 + d(n-1) → a25=-7+5(25-1)

OpenStudy (solomonzelman):

a25 = ?

OpenStudy (anonymous):

113

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

Now use the S(n) formula

OpenStudy (solomonzelman):

S(n)=½(a1+an)×n

OpenStudy (solomonzelman):

S(25) = ½ (-7 + 113) × 25

OpenStudy (anonymous):

so it would be 1325?

OpenStudy (solomonzelman):

Yes, very good !

OpenStudy (solomonzelman):

You are a pleasure to work with :)

OpenStudy (anonymous):

thank you so much for all you help

OpenStudy (anonymous):

all of your help* oops

OpenStudy (solomonzelman):

You welcome, have fun !